The sum of an angle plus its complement plus its supplement is 200 more than the angle. Find the measure of the supplement of the angle.
Answer by algebrahouse.com(1659)

"The sum of an angle plus its complement plus its supplement is 200 more than the angle. Find the measure of the supplement of the angle."

x = the angle
90 - x = its complement
180 - x = its supplement

x + 90 - x + 180 - x = x + 200 {sum of angle plus complement plus supplement is 200 more than angle}
-x + 270 = x + 200 {combined like terms}
-2x = -70 {subtracted 270 and subtracted x from both sides}
x = 35 {divided both sides by -2}
180 - x = 145 {substituted 35, in for x, into 180 - x}

145 is the supplement
www.algebrahouse.com